### v1

# Echo chamber effects in signed networks

Abstract
Echo chamber effects in social networks are generally attributed to the prevalence of interactions among like-minded peers. However, recent evidence has emphasized the role of hostile interactions between opposite-minded groups. Here, we model information propagation between such groups by generalizing popular contagion models to signed networks. We show that echo chambers spontaneously emerge in balanced networks, and in antibalanced ones for specific parameters. The robustness of our results is confirmed through simulations on various network topologies, including a real-world dataset.
Echo chambers, Signed networks, Structural balance, Social networks
The study of echo chambers has been central in research about online social platforms. By continuously reinforcing pre-existing views [1, 2], these tightly knit communities of like-minded users exacerbate the polarization of opinions and in turn, the fragmentation of societies [3, 4, 5, 6]. They have been observed at multiple levels, from the connections established by the users themselves, to personalised timelines of content curated by engagement-seeking algorithms [7, 8].
In recent years however, this evidence has been increasingly challenged [9, 10, 11, 12, 13, 14]. Scholars have pointed out that there is communication between opposite-minded communities, often in the form of hostile interactions [15, 16, 17]. Opposite-minded users may frame the same piece of information under different lights, reinforcing ideological divides while still allowing information to travel through partisan lines. Negative relations might be even more prevalent than positive ones for some users, such as journalists [16]. Thus, it is primordial that we distinguish between friends and foes when studying online debates [18].
The presence of hostile interactions among individuals is often modelled as a signed network with positive and negative links [19]. The inclusion of negative links gives rise to novel phenomena, such as structural balance [20, 21]. A signed network is perfectly balanced if it can be divided into two groups, so that all links between agents of the same group are positive and all links between agents of different groups are negative [20]. In other words, it contains two groups that are antagonistic towards one another [22, 23]. While perfectly balanced networks are a mathematical idealization, studies have consistently shown a tendency towards structural balance in empirical networked systems of various types [24, 25, 26, 27]. On the other hand, when all links within each group are mostly negative and links between groups are mostly positive, the network is said to be antibalanced [28]. While not as common as balanced networks, antibalanced networks can emerge in historical, social and political contexts, like international relations during the Thirty Years’ War [29] or the U.S. Congress during the Carter and Reagan administrations (see [22] and the Appendix, Sec. D).
These findings have motivated the introduction of the notion of degree of balance, with indices based on triangles [27, 22], random walks [30, 31, 32] or frustration metrics [33, 34], among others. Recent research has discovered intriguing connections between the degree of balance of social and political networks and their tendency towards polarization [22, 34]. The degree of balance also impacts the evolution of various dynamical processes [35, 36, 32], including contagion models [37, 38, 39], random walks [40], diffusive dynamics [41] and the voter model [42]. Conversely, structural balance can spontaneously arise as a result of opinion dynamics on signed networks [43, 44].
Within this context, our contribution is two-fold. First, we extend to signed networks two popular models of information propagation, the Independent Cascade Model and the Linear Threshold Model [45, 46]. The key idea driving these extensions is that a single piece of information may be framed in different ways between, and within, antagonistic groups. Then, we use those models to analyze the key role played by the degree of balance in the emergence of echo chambers. We uncover a sudden transition as the degree of balance becomes positive, and an anomalous region where echo chambers can emerge in networks with negative degree of balance. Echoing empirical observations on the correlation between activity and polarisation [47, 48, 1, 49], we find reinforced echo chambers when information spreads more easily. The robustness of our results is confirmed on different network topologies, including a real-life dataset. As a by-product, we also provide a simple procedure to transform an unsigned network into a signed one with a prescribed degree of balance.
Signed social networks.
We consider an undirected signed network with $N$ agents. Each agent is assigned a latent group membership, which represents the divergent opinions within the population. If all in-group interactions are friendly and all out-group interactions are hostile, then the network is perfectly balanced. In the opposite case, where in-groups are always hostile towards each other and friendly with out-groups, we call the network perfectly antibalanced [28]. To interpolate between these two extremes, we employ the triadic degree of balance $\tau$, commonly used in the literature [50, 27, 51, 22]. It is defined as:
$\tau=\text{tr}(W^{3})/\text{tr}(|W|^{3}),$ (1)
where $W$ is the signed adjacency matrix of the network and $|W|$ its entrywise absolute value. In perfectly balanced networks $\tau=+1$, and in perfectly antibalanced ones $\tau=-1$. Values in-between quantify intermediate degrees of balance, with $\tau\approx 0$ for networks that are neither significantly balanced nor antibalanced. The definition of $\tau$ is based on the product of link signs in triangles, which have been found to be an accurate and computationally fast proxy for balance  [^1] Triangles with a positive product of link signs famously respect the old adage “the friend of my friend is my friend, the enemy of my friend if my enemy, the enemy of my enemy is my friend”.. Hereafter, we employ the terms balanced and antibalanced to refer to networks with $\tau>0$ and $\tau<0$ respectively, while we reserve the terms perfectly (anti)balanced for the cases $\tau=\pm 1$.
The majority of our simulations are run on networks with an underlying Erdös-Rényi (ER) topology with connection probability $p$. Hence, they do not exhibit communities from an unsigned point of view. Nevertheless, group memberships can induce mesoscopic structures affecting the degree of balance. We thus attribute signs to the links following a procedure that guarantees a pre-specified triadic degree of balance on average. This procedure is inspired by [53] and detailed in the Appendix, Sec. A.
Information propagation.
We assume that an item is propagating throughout the network: a journal article, a tweet, a picture, etc. To describe its propagation, we adapt the archetypal Independent Cascade Model of simple contagion [45]. In this model, agents are initially all inactive (state 0), except for a few seeds who are the original emitters of the item (state 1). Their neighbors are thus exposed to the item, and may choose to share it themselves. Formally, in consecutive steps, each agent that was activated in the previous step activates each of its neighbors individually with probability $\lambda$. This probability is called the activation rate. If all attempted activations in a single step fail, the process stops. To account for signed links, we add a novel feature to these dynamics and propose the Signed Independent Cascade Model (SICM). Active agents can now be in one of two states $\{-1,+1\}$, reflecting two opposite-minded ways to frame the information: a fresh government decision may spark praise from some and criticism from others. When an agent is activated through a positive (resp. negative) link, they adopt the same (resp. opposite) state of the agent responsible for their activation. Once an agent is active, its state is fixed and never changes. More details can be found in the Appendix, Sec B.2.
The SICM relies on a simple contagion mechanism: one exposure to information is enough to trigger a potential activation. In contrast, complex contagion mechanisms necessitate multiple exposures. This creates a different phenomenology [54, 55, 56] that has been argued to better model realistic behavioral dynamics [57, 58, 59]. For this purpose, we propose the Signed Linear Threshold Model (SLTM), adapted from [45, 46]. Similarly as for the SICM, agents can be in one of two states $\{-1,+1\}$ depending on who triggered their activation. Agents activate if the aggregated influence of their neighbors exceeds a threshold $T$—modelling details can be found in the Appendix, Sec. B2. While most of our results are presented for the SICM, we show at the end that the SLTM exhibits very similar behaviors.
Echo chamber effect
To measure the intensity of the echo chamber effect, we adapt a metric from [60]. Let $\rho_{ab}^{+}$ denote the expected fraction of agents in state $+1$ among all the active agents in group $b$, at the end of a process with all seeds belonging to group $a$. Similarly, we define $\rho_{ab}^{-},\rho_{ba}^{+},\rho_{ba}^{-}$ as well as $\rho_{aa}^{+},\rho_{aa}^{-},\rho_{bb}^{+},\rho_{bb}^{-}$, excluding the seeds from the computation of the last four. The echo chamber strength is defined as
$E=\frac{\rho_{aa}^{+}-\rho_{aa}^{-}-\rho_{ab}^{+}+\rho_{ab}^{-}+\rho_{bb}^{+}-\rho_{bb}^{-}-\rho_{ba}^{+}+\rho_{ba}^{-}}{4}.$ (2)
In the Appendix, Sec. C.2. we prove that $E$ is a second-order moment; thus, we can interpret $E$ as a measure of the correlation between states and group memberships. The maximum $E=+1$ corresponds to perfect echo chambers, where everyone inside a group is in the same state, opposite to that of the other group. The minimum $E=-1$ is the case of an anti-echo chamber, where agents within the seed group disagree with the seeds, while agents of the opposite group agree with the seeds. When $E\approx 0$, there is no significant echo chamber effect. For the sake of conciseness, we refer to $E$ simply as the echo chamber effect, bearing in mind that it is actually a metric that quantifies this effect.
Results.
Figure 1: Echo chamber effect on signed ER networks of various densities $p$. On the right half of the heatmaps are balanced networks $(\tau>0)$, on the left half are antibalanced networks $(\tau<0)$.
Figure 2: Echo chamber effect in balanced ER networks. (a) $E$ as a function of the $\tau$ for various $\lambda$. (b) $E$ as a function of $\lambda$ for various $\tau$. (c) Activation rates $\lambda$ realising the minimum and maximum echo chamber effect for various $p$, averaged over all values of $\tau$. (d) Density $p$ realising the minimum and maximum echo chamber effect for various $\lambda$, averaged over all values of $\tau$. Shaded areas cover one standard deviation from the mean.
We assess the echo chamber effect via simulations of the SICM on ER networks with $N=250$ agents. Figure 1 shows $E$ as function of $\tau$ and $\lambda$ on networks of various densities $p$. The difference between balanced ($\tau>0$) and antibalanced ($\tau<0$) networks is immediately striking, with a transition between the two that gets sharper as $p$ increases. In balanced networks, $E$ increases monotonically with $\tau$, as shown in Fig.2(a). Thus, we find that a stronger balance entails a higher echo chamber effect. In other words, echo chambers, much like polarization [22, 31] and dissensus [41], are amplified by structural balance in signed social networks.
Furthermore, $E$ is in fact not monotonous with respect to $\lambda$, as shown in Fig. 2(b). There exists a nontrivial value of $\lambda$ for which $E$ is minimal, which decreases continuously with the network density (Fig. 2(c)). In contrast, the argmax of $E$ jumps discontinuously from $\lambda\approx 0$ to $\lambda\approx 1$. The same phenomenon is observed when computing the argmax of $E$ over $p$ (Fig.  2(d)). This suggests that minimizing the echo chamber effect is simpler than maximizing it: the optimal values of $\lambda$ and $p$ regarding minimization span a narrower range of values and are less susceptible to perturbations, whereas the optimal values for maximization vary widely between 00 and $1$ and can change abruptly in the critical region.
In antibalanced networks, one could expect the same behavior, with $E$ decreasing monotonously towards $-1$ as $\tau$ becomes more negative. As we see in Fig. 1 however, this is not the case. Instead, we find that $E$ changes non-monotonously with both $p$ and $\lambda$ and $\tau$ (see Appendix, Sec. D). Specifically, $E$ reaches its minimum value, $E=-1$, both when $p$ and $\lambda$ are very high (e.g. $p=\lambda=1$) and when they are very low (e.g. $p=0.025,\lambda=0.01$). Perhaps most surprisingly, certain regions of the $(\tau<0,\lambda)$ half-plane display positive echo chamber effects. This anomalous region, corresponding to the red semi-circular area of Fig. 1, shifts downwards as the network density $p$ increases, and almost vanishes at high densities. Interestingly, Figure 3(a) reveals that for any given $\tau$, the boundaries of the anomalous region with respect to $p$ and $\lambda$ do not move. Instead, the value of $\tau$ only influences the strength of this anomaly.
Figure 3: (a) Echo chamber effect as a function of network density and activation rate in antibalanced ER networks, for three different $\tau$. (b) Two realisations of the SICM on a signed ER network with $p=0.5,\tau=-1$, for $\lambda=0.08$ (echo chamber) and $\lambda=0.30$ (anti echo chamber). Circles correspond to agents of group $a$ and triangles to agents of group $b$. Agents are colored according to their state: grey for state 00, green for state $+1$ and pink for state $-1$.
To better understand this phenomenon, we compare how the propagation unfolds for two different values of the activation rate. Figure 3(b) shows simulations of the SICM on a highly dense, perfectly antibalanced network ($\tau=-1$) with $\lambda=0.08$ and $\lambda=0.30$. Because $\tau=-1$, agents activated during odd steps adopt state $-1$ if they are in the seed group and state $+1$ otherwise, while the opposite happens during even steps. For $\lambda=0.30$, most activations occur in the first step, leading the seeding group to predominantly adopt state $-1$, resulting in a negative echo chamber effect. However, with $\lambda=0.08$, fewer activations happen initially, and most agents are activated in the second step. This causes the seeding group to adopt state $+1$, and generates a positive echo chamber effect even though the network is perfectly antibalanced. This alternance mechanism leads us to believe that this phenomenon is due to the synchronous aspect of the propagation process. Indeed, we show in the Appendix, Sec. D that these anomalies disappear with asynchronous activations. Additionally, the exact shape of the anomalous regions seem to change slightly with the network size (Appendix, Sec. D).
Robustness.
To demonstrate the robustness of our approach, we show that similar results are achieved (i) on networks with different topologies, and (ii) with our complex contagion model. In Fig. 4, we apply the SICM on a collection of networks with similar densities but various topological features. The Stochastic Block Model (SBM) generates networks with community structure, the model of Watts-Strogatz (WS) generates networks with clustering, and the model of Barabasi-Albert (BA) generates scale-free networks. Additionally, we consider a network extracted from a real-life dataset and exhibiting all of the above, Polblogs [61] (details in the Appendix, Sec. D). The echo chamber effect emerges when $\tau>0$ for every network, confirming that structural balance favors echo chambers regardless of the topology. Moreover, the anomalous region is also present for every network except the Watts-Strogatz one (Fig. 4f), due to clustering hindering the dominance of even-length paths.
For complex contagion, we rely on the SLTM. Agents can now only activate if a sufficiently high number of their neighbors are active themselves. Even with this fundamental difference, our main results regarding the emergence of echo chambers hold. Figure 5 shows that SLTM dynamics again lead to a positive echo chamber effect when $\tau>0$, and an anomalous region in the $\tau<0$ half-plane. Slight differences with the SICM include are the sharper transitions, an increased presence of the anti-echo chamber effect in the $\tau<0$ half-plane, and the persistence of a few anomalous regions under asynchronous activations (Appendix, Sec. D). Overall, these results suggest that the echo chamber phenomenology is robust with regards to the dynamical model employed.
Figure 4: Echo chamber effect with different networks topologies. (a) ER network with $p=0.025$. (b) SBM network with in-group link density $p_{in}=0.04$ and out-group link density $p_{out}=0.01$. Community memberships match group memberships. (c) WS network with average degree $k=6$ and rewiring probability $q=0.5$ (low clustering). (d) BA network with average degree $k=6$. (e) Polblogs network [61]. (f) WS network with average degree $k=6$ and rewiring probability $q=0.1$ (high clustering). All networks have and $N=250$ agents, except Polblogs with $N=1,222$.
Figure 5: echo chamber effect resulting from a complex contagion process (SLTM) across various topologies. Same networks as in Fig. 4.
Discussion.
Our study makes two key contributions. First, we propose two opinion contagion models, the SICM and the SLTM, that address how information can be framed differently among antagonistic groups. Second, and more importantly, we employ these models to describe the role of structural balance and connectivity in the emergence of echo chambers. Our findings align with previous results connecting polarization with structural balance [31, 22, 34] and connectivity [43]. However, while those studies focused on structural properties, ours is the first to emphasize the impact of these on information spreading. Our results hold for both simple and complex contagion and in a variety of network topologies, as well as in real networks.
This work represents a first step towards incorporating the effect of negative links in the study of echo chambers, with potential extensions to incorporate temporal activity of agents [1], non-Markovian dynamics [62] or homophily [60, 63]. Moreover, the SICM and SLTM show nontrivial regimes and transitions that could be explored more thoroughly. Future works could also consider the effects of structural balance beyond triangles [51, 30], other information transmission biases [60] and examine echo chambers at microscopic and mesoscopic levels [51]. The comparison between theoretical predictions and empirical data (e.g. [6, 57]) is also key to assess the validity of our modelling assumptions. Finally, our results could be applied to mitigate echo chamber effects and polarization in recommendation algorithms [64].
Acknowledgements.
We want to thank the young researchers of the Complex Systems Society for financing a research visit of AV to IFISC through the Bridge Grants program, that served as the primary basis for this work. We also thank Maxi San Miguel for insightful discussions and David Abella for suggesting improvements to the first draft. FDD thanks financial support from MDM-2017-0711-20-2 funded by MCIN/AEI/ 10.13039/501100011033 and FSE, as well as project APASOS (No. PID2021-122256NB-C22). In the early stages of this research, AV was funded by the UK EPSRC grant EP/S022503/1 that supports the Centre for Doctoral Training in Cybersecurity delivered by UCL’s Departments of Computer Science, Security and Crime Science, and Science, Technology, Engineering and Public Policy.
Appendix A Networks generation
To generate signed networks in which to run the dynamical processes, we always follow a three-steps procedure. First, we create an unsigned network through standard generative algorithms, like the Erdös-Rényi or Barabasi-Albert ones. Second, we assign a group to each of the agents in the network. In Erdös-Rényi, Watts-Strogatz and Barabasi-Albert networks, each agent is assigned randomly to one of two groups of the same size, whereas for the Stochastic Block Model and Polblogs, the group labels correspond to the communities existing in the network. Third, we attribute signs to links as follows: links whose endpoints fall within the same group are negative with probability $\eta$ and positive otherwise; whereas links whose endpoints fall in different groups are positive with probability $\eta$ and negative otherwise. The parameter $\eta$ regulates the degree of balance in the network. For $\eta=0$ the network is (perfectly) balanced, for $\eta=1/2$ link signs are uncorrelated with group memberships, while for $\eta=1$ we obtain an antibalanced network. Intermediate values allow us to interpolate between the different degrees of balance. One may then think of $\eta$ as a noise parameter.
To generate networks with a specific triadic degree of balance $\tau$, we make use of the following nonlinear relation, proven in Sec. C.1 below:
${\tau}=(1-2\eta)^{3}$ (3)
on expectation. Therefore, setting $\eta=(1-\tau^{1/3})/2$ will yield an ensemble of networks with average degree of balance $\tau$. This result is topology-independent and holds for any given network. We use it to control precisely the degree of balance of the networks we generate for our experiments. The relation is illustrated in Fig. 6 for networks of various topologies.
Figure 6: Degree of balance as a function of noise, for various networks: ER with density $p=0.1$, ER with density $p=1.0$, BA with average degree $k=3$, and Polblogs [61]. Each point is averaged over 10 random assignments of link signs. Vertical bars indicate one standard deviation from the mean.
Polblogs is a real-life dataset describing connections between political blogs in the US in 2003 [61]. Each node corresponds to a blog and is in one of two communities, Democrat or Republican. The presence of a link between two agents indicates that one mentioned the other. We set group memberships to correspond to community memberships. We restrict ourselves to the largest connected component of the network, which has $1,222$ agents and $19,021$ links.
Table 1 summarizes the main characteristics of the networks. We highlight the fact that Polblogs exhibits both a community structure as the SBM, and local clustering as the WS model. While it is not exactly scale-free like the BA model, its degree distribution is well-fitted by a power-law with exponential cutoff (see [61]), which implies the presence of few hubs and many leaves.
ER BA WS $(q=0.1)$ WS $(q=0.5)$ SBM Polblogs
Average density 0.025 0.024 0.024 0.024 0.025 0.0224
Hubs and leaves - - - -
Communities - - - -
Average modularity - - - - 0.84 0.41
Average clustering coefficient 0.02 0.08 0.45 0.1 0.03 0.32
Table 1: Basic statistics of the networks considered.
Appendix B Simulation details
B.1 Signed Independent Cascade Model (SICM)
The simulation process is as follows. To compute $E$ for Erdös-Rényi networks with density $p$ and degree of balance $\tau$, we first generate $m=20$ such networks via the process described in Sect. A above. Then we perform $2m$ of the SICM with activation rate $\lambda$ on each of these networks, with $m$ simulations seeded in group $a$, and $m$ simulations seeded in group $b$. The number of seeds is chosen to be $10\%$ of the seeding group every time. The remaining agents are initially inactive. At each step, each agent that was activated in the previous step tries to activate independently each of its neighbours. The probability that a given neighbor is activated this way, is the activation rate $\lambda$. If an agent becomes activated through a positive link, it inherits the state ($+1$ or $-1$) of the agent that activated it. If it is activated through a negative link, it instead adopts the opposite state. Once an agent has tried to activate all its inactive neighbors, it ceases to try to activate them in future steps—in other words, each agent tries to activate its neighbors a single time.
For each simulation, we use a breadth-first search algorithm. Contrary to the traditional Independent Cascade, here the specific agent that activates another one matters in the SICM. For example, consider a inactive agent with two neighbors in state $+1$, where one of the neighbors is connected through a positive link and the other through a negative one. The final state of the agent will change depending on which of the agents activates it. Moreover, since the agent cannot update its state once activated, the fist agent that tries to activate has a better chance of propagating its state. To circumvent this issue, we randomize the order in which neighbors try to activate the agent. Finally, we restart the simulation if no agents were activated in a single community. This way, the values of $\rho$ and $E$ are well defined.
We also compare the results with an asynchronous version the process. Whenever an agent becomes activated, each of its neighbors are placed in a queue. Then at each step, an inactive agent is picked at random from the queue and activates with probability $\lambda$. If it succeeds, it adopts the same state as the agent that triggered this activation. This has a clear impact on the results, as the anomalous regimes in the $\tau<0$ region disappear (Figs. 10, 11).
B.2 Signed Linear Threshold Model (SLTM)
We propose the SLTM to adapt the archetypal Threshold Model [45] for signed networks. As for the SICM, all agents are initially in state 00, except a few seeds in state $+1$, who all belong to the same group. At each step, every inactive agent $i$ is selected. The average influence from its neighbours is computed as:
$\phi_{i}(t)=\frac{1}{k_{i}}\sum_{j=1}^{k_{i}}w_{ij}x_{j},$ (4)
where $k_{i}$ is the degree of agent $i$ (independent of the signs of the links), $x_{j}\in\{-1,0,+1\}$ is the state of agent $j$ and $w_{ij}$ the sign of the link between $i$ and $j$. Notice that both $w_{ij}$ and $x_{j}$ can be negative, so $\phi_{i}\in[-1,1].$ We then update the state of $i$ as follows.
$x_{i}\leftarrow\begin{cases}+1&\text{if}\quad\phi_{i}(t)>T,\\ +1&\text{if}\quad\phi_{i}(t)=T\quad\text{with probability }1/2,\\ -1&\text{if}\quad\phi_{i}(t)<-T,\\ -1&\text{if}\quad\phi_{i}(t)=-T\quad\text{with probability }1/2,\\ 0&\text{otherwise}.\end{cases}n$ (5)
The threshold $T\in(0,1)$ quantifies the difficulty with which information flows through the network. Thus, $1-T$ is akin to $\lambda$ in the SICM, as it measures how easily information travels. Note that all activations happen simultaneously each step. In Fig. 12 we compare the results with an asynchronous version the process, where agents are activated one by one in a random order. This has a clear impact on the results, as the anomalous regimes in the $\tau<0$ region (almost) disappear.
The simulation process is similar as for the SICM. To compute $E$, we first generate $m=20$ such networks via the process described in Sect. A above. Then we perform $2m$ of the SLTM with threshold $T$ on each of these networks, with $m$ simulations seeded in group $a$, and $m$ simulations seeded in group $b$. The number of seeds is chosen to be $10\%$ of the seeding group every time.
Appendix C Mathematical proofs
C.1 Triadic balance as a function of noise
We prove that for any network,
$\mathbb{E}[\tau]=(1-2\eta)^{3},$ (6)
where $\tau$ is the degree of triadic balance and $\eta$ is the level of noise as defined in Sec. A above. Given a signed network, consider a triangle $T$ in this network. The triangle has its own degree of balance $\tau(T)$. In the absence of noise ($\eta=0$), we have $\tau(T)=1$. If $\eta>0$, each link in this perfectly balanced version of $T$ is flipped with probability $\eta$.
  • With probability $(1-\eta)^{3}$ no link is flipped, and $T$ stays balanced.
  • With probability $3(1-\eta)^{2}\eta$ exactly one link is flipped, and $T$ becomes antibalanced.
  • With probability $3(1-\eta)\eta^{2}$ two links are flipped, and $T$ stays balanced.
  • With probability: $\eta^{3}$ all three links are flipped, and $T$ becomes antibalanced.
The expected degree of balance of $T$ after the flips is
$\mathbb{E}[\tau(T)]$ $=(1-\eta)^{3}-3(1-\eta)^{2}\eta+3(1-\eta)\eta^{2}-\eta^{3}$
$=(1-2\eta)^{3}.$ (7)
The expected degree of balance $\mathbb{E}[\tau]$ of the whole graph is the average of $\mathbb{E}[\tau(T)]$ over all triangles $T$ after the flips, hence $\mathbb{E}[\tau]=(1-2\eta)^{3}$ as well.
C.2 The echo chamber effect is a second order moment
Let $a,b$ denote two complementary subsets of $V$ representing the two groups in the network, and let $f$ be a vector indicating the group of each agent: $f_{i}=+1$ if $i\in a$, $f_{i}=-1$ if $i\in b$. Similarly, let $I^{\pm}$ represent the set of agents with positive and negative state at the end of the contagion process, and let $x$ be a vector defined as follows: $x_{i}=\pm 1$ if $i\in I^{\pm}$, $x_{i}=0$ otherwise. Furthermore, let $s$ denote the set containing the seeds. Based on these definitions, we can express the fraction of agents in a final state ±$\pm 1$, conditioned on the seeds belonging to group $F_{s}$ and the analyzed agents being active and belonging to group $F_{t}$, as:
$\rho_{F_{s}F_{t}}^{\pm}=\frac{|\left\{i\in F_{t}:x_{i}=\pm 1,s\subset F_{s}\right\}|}{N_{I}(F_{s},F_{t})}$ (8)
where $N_{I}(F_{s},F_{t})$ denotes the final number of active agents belonging to group $F_{t}$. Note that this quantity also depends on the seed group $F_{s}$, as the seed group can affect the efficiency of the contagion process and thus the final number of active agents.
By the law of large numbers, and assuming independent final states (pair approximation [65]), the fraction of active agents in a final state $x_{i}\in\{\pm 1\}$ is approximately equal to the probability of a final state $x_{i}$ in agent $i$ conditioned on $i$ being active and belonging to a given target group $F_{t}$, and the set of seeds belonging to group $F_{s}$:
$\rho_{F_{s}F_{t}}^{\pm}$ =|$=\frac{|\left\{i\in F_{t}:x_{i}=\pm 1,s\subset F_{s}\right\}|}{N_{I}(F_{s},F_{t})}$
$\approx\mathbb{P}[x_{i}=\pm 1|x_{i}\neq 0,s\subset F_{s},i\in F_{t}].$ (9)
To obtain an unconditioned expectation of the number of active agents on a given state, we employ the law of total expectation, $\mathbb{E}_{X,Y}[X]=\mathbb{E}_{Y}[\mathbb{E}_{X}[X|Y]]$, where the subindex indicates the random variable over which the expectation is taken. In this case, the conditional random variable $Y$ corresponds to both the seed and target groups $F_{s}$ and $F_{t}$:
$\mathbb{E}_{x,F_{s},F_{t}}[x_{i}|x_{i}\neq 0]$ $=\mathbb{E}_{F_{s}}[\mathbb{E}_{F_{t}}[\mathbb{E}_{x}[x_{i}|x_{i}\neq 0,s\subset F_{s},i\in F_{t}]]]$
$=\frac{1}{4}\sum_{F_{s}\in\{a,b\}}\sum_{F_{t}\in\{a,b\}}\sum_{x_{i}\in\{\pm 1\}}x_{i}\mathbb{P}[x_{i}|x_{i}\neq 0,s\subset F_{s},i\in F_{t}]$
$=\frac{1}{4}\sum_{F_{s}\in\{a,b\}}\sum_{F_{t}\in\{a,b\}}\sum_{x_{i}\in\{\pm 1\}}x_{i}\rho^{x_{i}}_{F_{s}F_{t}}$
$=\frac{1}{4}\left(\rho^{+}_{aa}-\rho^{-}_{aa}+\rho^{+}_{ba}-\rho^{-}_{ba}+\rho^{+}_{ab}-\rho^{-}_{ab}+\rho^{+}_{bb}-\rho^{-}_{bb}\right).$ (10)
where the factor $1/4$ comes from assuming equally likely group membership, so that
$\mathbb{P}(s\subset a)=\mathbb{P}(s\subset b)=\mathbb{P}(i\in a)=\mathbb{P}(i\in b)=1/2.$ (11)
We have thus proved that the expected value of $x_{i}$ (conditioned on being active) is linearly dependent on the variables $\rho_{ab}^{\pm}$, accounting for the sign. We can now do a similar argument to relate the echo chamber effect to these variables. To do so, we first define a variable $c_{si}$ indicating whether the seed and target groups agree:
$c_{si}=\begin{cases}+1\quad\textrm{if}\quad f_{s}=f_{i}\\ -1\quad\textrm{if}\quad f_{s}\neq f_{i},\end{cases}$ (12)
where $i$ is an arbitrary agent, $s$ is the set of seeds and all seeds belong to the same group. Alternatively, we can express $c_{si}$ as a product of the groups of $i$ and $s$: $c_{si}=f_{s}f_{i}$. Note that, since the group assignations can be random, $c_{si}$ is a random variable and we may define a joint probability of $x_{i}$ and $c_{si}$. We now prove that $E$ is a second-order moment of this joint distribution using the methodology employed before:
$\mathbb{E}_{x,F_{s},F_{t}}[x_{i}c_{si}|x_{i}\neq 0]=$
$=\mathbb{E}_{F_{s}}[\mathbb{E}_{F_{t}}[\mathbb{E}_{x}[x_{i}c_{i}|x_{i}\neq 0,s\subset F_{s},i\in F_{t}]]]$
$=\frac{1}{4}\sum_{F_{s}\in\{a,b\}}\sum_{F_{t}\in\{a,b\}}\sum_{x_{i}\in\{\pm 1\}}x_{i}f_{i}f_{s}\times$ (13)
×$\quad\quad\quad\times\mathbb{P}[x_{i}|x_{i}\neq 0,s\subset F_{s},i\in F_{t}]$
$=\frac{1}{4}\left(\rho^{+}_{aa}-\rho^{-}_{aa}-\rho^{+}_{ba}+\rho^{-}_{ba}-\rho^{+}_{ab}+\rho^{-}_{ab}+\rho^{+}_{bb}-\rho^{-}_{bb}\right)$ (14)
The last expression is the echo chamber effect, so we have proven that $E$ is a second-order moment of the joint distribution of the final states $x_{i}$ and the seed-target group agreeement variable $c_{si}$.
Appendix D Additional figures
Figure 7 shows an example of antibalance in a real-world, signed network from [22]. The agents are members of the U.S. congress and the sign of links connecting them is based on bill co-sponsorships. We see that during the Carter and Reagan administrations (1977-1985), the House of Representatives exhibits a negative degree of balance, and so does the Senate between 1979 and 1983 (96th-97th congress). Thus, in these instances, congressmembers were co-sponsoring bills with colleagues from the other party more than within their own ranks.
Figure 8 gives another perspective on the highly heterogeneous evolution of $E$ with $\lambda$ and $\tau$ in antibalanced networks and under the SICM. Also for antibalanced networks and SICM, Fig. 9 illustrates how the anomalous region changes for simple contagion with a different value of $N$, here $100$ instead of $250$ in the article. The anomalous region has a highly similar shape but is slightly bigger. Finally, Figs. 10 and 11 show that the anomalous region disappear with asynchronous updates in the SICM. They also mostly disappear in the SLTM, except for a tiny one that persists in the Polblogs network (Fig. 12). This last plot was obtained using the SLTM—see Sec. B.2 above for more details.
Figure 7: Degree of Balance in both chambers of the U.S. congress, from 1973 (93rd congress) to 2017 (111st congress).
Figure 8: Echo chamber effect in antibalanced ER networks. (a) $E$ as a function of the $\tau$ for various $\lambda$. (b) $E$ as a function of $\lambda$ for various $\tau$.
Figure 9: Echo chamber effect as a function of network density and activation rate in antibalanced ER networks with $N=100$ agents, for three different $\tau$.
Figure 10: Echo chamber effect in the SICM with asynchronous updates, on ER networks with various densities $p$.
Figure 11: Echo chamber effect in the SICM with asynchronous updates. (a) ER network with $p=0.025$. (b) SBM network with in-group link density $p_{in}=0.4$ and out-group link density $p_{out}=0.1$. Community memberships match group memberships. (c) WS network with average degree $k=6$ and rewiring probability $q=0.5$ (low clustering). (d) BA network with average degree $k=3$. (e) Polblogs network [61]. (f) WS network with average degree $k=6$ and rewiring probability $q=0.1$ (high clustering). All networks have and a density between $0.022$ and $0.025$, and $N=250$ agents—except Polblogs with $N=1,222$.
Figure 12: Echo chamber effect in the SLTM with asynchronous updates. (a) ER network with $p=0.025$. (b) SBM network with in-group link density $p_{in}=0.4$ and out-group link density $p_{out}=0.1$. Community memberships match group memberships. (c) WS network with average degree $k=6$ and rewiring probability $q=0.5$ (low clustering). (d) BA network with average degree $k=3$. (e) Polblogs network [61]. (f) WS network with average degree $k=6$ and rewiring probability $q=0.1$ (high clustering). All networks have and a density between $0.022$ and $0.025$, and $N=250$ agents—except Polblogs with $N=1,222$.